Given =2x2-9x+2, =1-6x and =x2-4, find each function. Indicate any restrictions in the domain.

Given =f(x)2x2-9x+2, g(x) =1-6x and h(x) =x2-4, find each function. Indicate any restrictions in the domain.

(f\cdot h) (x)
Answer #1

solution dau : 242 -9xt 2 gay = 1-605 how = x²-4 Ifahy z _ farh602 (4-90+2) (22²-4) se? - 9 x3 +2 2² 16+36x -8 -9x3 +6x² + 36